### v1

# The Surprising Effectiveness of Test-Time Training for Abstract Reasoning

Abstract
[…]
1 Introduction
Large-scale neural language models (LMs) excel at performing tasks that occur in their training data, and often elementary variations or compositions of those tasks (Brown et al., 2020; Todd et al., 2024). Given natural language task specifications or a small number of examples, LMs often successfully infer the desired task and produce an appropriate output. But can LMs also solve new problems, involving non-trivial reasoning, planning, or string manipulation of a kind very different from their pre-training data? This question is central to understanding the novel skill acquisition capabilities of current AI systems, which has been proposed as a key measure of intelligence (Chollet, 2019).
For complex and novel tasks, it is often difficult to obtain a correct answer simply by sampling from an LM (Wu et al., 2023). However, a significant finding in recent years has been that LM performance can be substantially improved by augmenting LM decoding with additional test-time computation. Methods in this category include chain-of-thought prompting (Wei et al., 2022), sampling with majority voting (self-consistency; Wang et al., 2022), code execution (Brown et al., 2024; Snell et al., 2024; Damani et al., 2024), and search (Yao et al., 2024).
One scaling strategy that has gained recent attention is test-time training (TTT), in which models are updated through explicit gradient steps based on test-time inputs (Krause et al., 2018; 2019). This method differs from standard fine-tuning as it operates in an extremely low-data regime—typically via an unsupervised objective on a single input, or a supervised objective applied to one or two in-context labeled examples. Modern versions of this approach was proposed for vision models by Sun et al. (2020), and also applied to sequence models by Gandelsman et al. (2022). The design space for TTT approaches is large, and there is currently a limited understanding of which design choices are most effective for LMs (and specifically for novel-task learning). In this paper, we systematically study the impact of various TTT design choices, as well as its interaction with pre-training and sampling schemes.
We evaluate these methods in the Abstraction and Reasoning Corpus (ARC) (Chollet, 2019), a collection of extremely challenging few-shot visual reasoning problems. ARC is an ideal benchmark for testing the limits of LM generalization as it presents novel tasks, in a novel format, requiring nontrivial search and inference capabilities. Current language models perform poorly on ARC. Most successful approaches have relied on program synthesis techniques (Butt et al., 2024; Ainooson et al., 2023; Huang et al., 2023), though recently Cole et al. (2024) reported promising results using TTT on the benchmark.
We identify several crucial ingredients for effective application of TTT to few-shot learning: (1) initial fine-tuning on synthetic tasks similar to those encountered at test time, (2) an augmented, leave-one-out task generation strategy for constructing the test-time dataset, (3) per-instance adapter training and (4) a self-consistency (Wang et al., 2022) approach under invertible transformations. With careful choices of these components, TTT can significantly improve LM performance on ARC—increasing accuracy by up to a factor of six over a 1B model, and achieving state-of-the-art results for published, purely neural models on the ARC task with a 8B model. Indeed, our results show that when equipped with test-time training, ordinary LMs can match or exceed the performance of many neuro-symbolic approaches on ARC.
Our main contributions [^1] Our implementation can be found at this link. are:
  1. 1.
    We identify and systematically analyze the key components needed for test-time training on ARC tasks, with a a novel test time training data generation and self-consistency component.
  2. 2.
    We achieve state-of-the-art results among published neural approaches on the ARC validation set:
    • 53% accuracy on the public validation set with a 8B parameter model.
    • 61.87561.875% accuracy when ensembled with program synthesis approaches, comparable to average human performance on the dataset.
  3. 3.
    We demonstrate that tasks that could only be solved by program synthesis previously can be solved with fully neural approaches equipped with our TTT framework.
These results challenge the assumption that symbolic components are strictly necessary for solving such complex tasks. Instead, they suggest that the critical factor in solving novel reasoning problems may be the allocation of proper computational resources during test time, perhaps independently of whether these resources are deployed through symbolic or neural mechanisms.
2 Preliminaries
In this section, we first formally describe the ARC challenge. Next, we give an overview of in-context learning and test-time training, which form the foundation of our investigation. Finally, we detail our default experimental setup.
2.1 ARC Challenge
The Abstraction and Reasoning Corpus (ARC) aims to evaluate the abstract reasoning capabilities of language models through their ability to solve visual puzzles. Each puzzle, henceforth referred to as task, is comprised of input-output pairs of 2-D grids (up to $30\times 30$ in size) that contain shapes or patterns made with up to 10 different colors, as displayed in Fig.˜1(b). The output of each pair is obtained by applying an intuitive and shared transformation rule or function $y=f(x)$. In practice, these transformations are highly diverse and composite, ranging from simple concepts such as reflection and counting, to more complex ones such as application of gravity and path finding.
Each task in ARC is composed of a training and test split, with:
  • Training examples denoted $(x^{\textrm{train}}_{k},y^{\textrm{train}}_{k})_{k=1}^{K}$ (typically $K$ ranges from 2 to 7).
  • Test examples denoted $(x_{m}^{\textrm{test}},y_{m}^{\textrm{test}})_{m=1}^{M}$ (typically $M$ ranges from 1 to 3).
Given the set of training examples, the goal is to predict the test output $y^{\textrm{test}}$ for test test input $x^{\textrm{test}}$ by reasoning about the underlying transformation.
We denote a task as $d=(\mathbf{x}^{\textrm{train}},\mathbf{y}^{\textrm{train}},\mathbf{x}^{\textrm{test}},\mathbf{y}^{\textrm{test}})$ where $d\in\mathcal{D}_{\textrm{ARC}}$, the collection of such ARC tasks. The original training and validation sets of ARC dataset, respectively $\mathcal{D}^{\textrm{train}}_{\textrm{ARC}}$ and $\mathcal{D}^{\textrm{val}}_{\textrm{ARC}}$, consists of 400 tasks each. Success criteria requires to produce exact match for all test outputs (if not partial points are given). Please refer to Johnson et al. (2021) for a taxonomy and analysis of these tasks.
Most approaches to ARC can be categorized into two main categories: program synthesis and fully neural. Program synthesis approaches (Butt et al., 2024; Wang et al., 2024; Li et al., 2024; Greenblatt, 2024) try to first find the transformation function $f$, and later apply it to the test example. On the other hand, fully neural approaches (Thoms et al., 2023; Bober-Irizar and Banerjee, 2024) try to directly predict the output $y^{\textrm{test}}$, only implicitly reasoning about the underlying transformation. In this work, we use a fully neural approach, using a LM to predict the test outputs.
We start with an LM pre-trained on text data (without a vision encoder). To provide ARC examples as input to these models, we thus require a formatting function (denoted str) that converts 2D grids into their textual representations as shown in Fig.˜8. Previous work has presented examples as lists of numbers (Wang et al., 2024) or color words, or lists of connected components labeled with shapes and locations (Greenblatt, 2024). Given any such string representation of a task, we may present it to an LM and perform predictions with few-short prompting, as explained in the next section.
2.2 In-context Learning
At a certain scale, many LMs exhibit the ability to adapt to new tasks without updating their parameters by simply conditioning on input examples or instructions provided. Given a sequence of input-output pairs $(x_{1},y_{1}),...,(x_{n},y_{n})$ and a new input $x_{n+1}$, a LM can be used to generate the output $\hat{y}_{n+1}$ by sampling from:
$\hat{y}_{n+1}\sim\textrm{LM}(\cdot\mid x_{1},y_{1},\dots x_{n},y_{n},x_{n+1})$ (1)
The possibility of in-context learning as implicit machine learning simulation discussed in previous work (Akyürek et al., 2022), but the empirical evidence shows that in-context learning with language models does not always resemble any standard machine learning algorithm (Zhao et al., 2024; Min et al., 2022), and it does not always work out-of-the box for novel tasks — e.g. small language models (few billion parameters) performs poorly on ARC (Opielka et al., 2024; Bober-Irizar and Banerjee, 2024).
2.3 Test-Time Training
Figure 2: TTT dataset generation for a test task (Section˜3.1): We start by creating leave-one-out tasks from the given training examples of the task. These tasks are then augmented through rule-based transformations to obtain the full TTT dataset. Finally, we train task-specific LoRA adapters on top of the base FT model.
Test-time training (TTT) enables parametric models to adapt during inference through dynamic parameter updates, an approach that remains relatively unexplored in the era of large language models. This technique is a form of transductive learning, where models leverages the test data structure to improve its predictions. The general TTT process works as follows: Starting with initial model parameters ${\bm{\theta}}_{0}$, for each test input (or batch of inputs), we first generate training data $\mathcal{D}_{\textrm{TTT}}(d_{\textrm{input}})$ from the test inputs. We then optimize these parameters to minimize a loss function $\mathcal{L}(\mathcal{D}_{\textrm{TTT}};{\bm{\theta}})$, producing temporarily updated parameters ${\bm{\theta}}_{d}$ for prediction. After generating predictions, the model is restored to the original parameters ${\bm{\theta}}_{0}$ for the next instance or batch. Thus, TTT trains a specialized prediction model for each test input, obtained by fine-tuning a base model on a test-time dataset generated from that test input.
In past work (e.g. Sun et al., 2020), $\mathcal{D}_{\text{TTT}}$ is typically constructed by applying an unsupervised objective (e.g. masked autoencoding) to the input $\mathbf{x}$ alone. However, the in-context learning setting we consider provides richer context in the form of demonstration pairs $(x_{1},y_{1}),\ldots,(x_{K},y_{K})$. Here, applying test-time tuning involves first constructing an initial language model $\mathrm{LM}$, mapping each test input $x$ to an input-specific dataset $\mathcal{D}_{\mathrm{TTT}}$, fine-tuning the LM to optimize some loss function $\mathcal{L}$ over the dataset according to: $\sum_{d\in\mathcal{D}_{\mathrm{TTT}}}\mathcal{L}(\mathrm{LM}(d))$, and finally sampling from the updated model to obtain a final prediction. Our experiments in this paper characterize each component of this pipeline, describing:
  1. 1.
    How to construct the augmented TTT dataset $\mathcal{D}_{\text{TTT}}$ from the test input (Section˜3).
  2. 2.
    An augmented inference strategy based on self-consistency over transformations (Section˜4).
  3. 3.
    A base model with parameters ${\bm{\theta}}_{0}$ that is fine-tuned on a dataset $\mathcal{D}_{\text{FT}}$ of similar tasks (Section˜5).
2.4 Experimental Setup
To investigate the impact of each TTT component, we conduct experiments by varying one component while holding the others constant at their optimal values (described in their respective sections). Our default configuration in the experiments uses the following settings:
Model Architecture & Optimization
We use an 8B parameter language model from the Llama-3 models, and 1B, 3B from Llama-3.2 models (Dubey et al., 2024). We use Low-Rank Adaptation (LoRA) (Hu et al., 2021) for parameter-efficient test-time training. For each task $d$, we initialize a separate set of LoRA parameters that are trained on the dataset $\mathcal{D}_{\textrm{TTT}}$. The LoRA rank is set to 128, and adaptations are applied to MLP, attention, and output layers. We train models with AdamW optimizer (Loshchilov and Hutter, 2019) with 2 epochs with batch sizes of 2.
Data & Formatting
For efficient evaluation purposes, we randomly pick 80 balanced ARC tasks from ARC validation set, includes 20 easy, 20 medium, 20 hard, 20 expert tasks according to the classification in LeGris et al. (2024a) (see Table˜2 for this task list). We will use this subset of ARC tasks throughout the paper, except our final results given in for the full validation set (Table˜1). We limit $\mathcal{D}_{\textrm{TTT}}$ to have maximum of 250 examples per task for efficiency reasons. With that, the whole TTT and inference process takes approximately 12 hours for 100 randomly sampled validation tasks when using an NVIDIA-A100 GPU. Section˜B.2 provides additional details on the hyper-parameters. Input grids are converted to text using numpy’s default array printing format as shown in Fig.˜8.
In the following sections, we investigate the key factors that contribute to successful abstract reasoning with language models. Our analysis covers the impact of fine-tuning data $\mathcal{D}_{\textrm{FT}}$ data, TTT data $\mathcal{D}_{\textrm{TTT}}$, training objectives, inference procedures, and model size, providing insights into effective strategy for deploying test-time training.
3 What Dataset and Loss During TTT?
3.1 Data Generation
Given a task, we take the set of training input-output pairs $(x^{\textrm{train}}_{k},y^{\textrm{train}}_{k})_{k=1}^{K}$ and turn them into an augmented set of test-time-training tasks $\mathcal{D}_{\text{TTT}}$. We obtain $\mathcal{D}_{\text{TTT}}$ using a two-step process: First, we create a set of leave-one-out in-context learning tasks from the given training input-output pairs. Second, we use invertible rule-based transformations on this set to obtain an augmented dataset. This process is summarized in Fig.˜2.
Step 1 - Leave-one-out Tasks: By excluding the $j^{\textrm{th}}$ example pair from the training examples, we can create the following synthetic task:
$d_{j}^{\textrm{ICL}}=(\underbrace{(x_{k},y_{k})_{k\in\{1,...,K\}\setminus\{j\}}}_{\textrm{synthetic training examples}},\underbrace{{x_{j},\quad\quad y_{j}}}_{\textrm{synthetic test example}})\quad\text{where }j\in[1,K]$ (2)
where $d_{j}$ synthetic training task with the $j$-th example pair treated as the test case. We can generate $n$ different tasks, each containing n1n-1 example pairs. We further include two randomly permuted version of $d_{j}$ where we permute the order of the training examples.
Step 2 - Rule-based transformations: Consider an invertible transformation $t$ such that $t^{-1}(t(x))=x$. For every task obtained in step 1, we can use $t$ to generate a new augmented task $t(d_{j}^{\textrm{ICL}})$, where $t$ is applied to each individual grid in the task.
We choose simple transformations that preserve the fundamental relationships while introducing controlled variations such as rotation, flips, color permutation, example permutaton, size scaling, etc. The list and the description of these transformations are provided in Table˜3. Finally, we obtain
$\mathcal{D}_{\textrm{TTT-ICL}}=\{t(d_{j}^{\textrm{ICL}})\}\quad\text{for all $t,j$ pairs.}$ (3)
Baseline: End-to-End Learning Tasks
For comparison to the “test-time in-context learning” approach described above, we also evaluate an “test-time end-to-end learning” approach. We create a supervised dataset directly from the example demonstrations by treating each input-output pair as an independent training instance. Unlike the in-context learning setup, no context is used for prediction:
$d_{j}^{\textrm{E2E}}=(x_{j},y_{j})\quad\text{where }j\in[1,K]$ (4)
Note that this would be equivalent to leave-(n1n-1)-out task set in ICL setting as no training examples are provided as context. Similar to ICL case, we can apply rule-based transformations to augment the dataset:
$\mathcal{D}_{\textrm{TTT-E2E}}=\{t(d_{j}^{\textrm{E2E}})\}\quad\text{for all $t,j$ pairs.}$ (5)
This approach is computationally more efficient as it directly learns the input-output mapping without the overhead of managing demonstration context i.e. the few-shot prompt.
3.2 Optimization Objective
During test-time training, we optimize a set of task-specific parameters using LoRA (low-rank adaptation; Hu et al. (2021)) while keeping most of the base model frozen. This approach allows computationally efficient adaptation while maintaining the model’s general capabilities.
Training Objective:
Given a task’s test-time training dataset $\mathcal{D}^{i}_{\textrm{TTT}}$, we minimize the standard language modeling loss on both the demonstrations and test outputs:
$\mathcal{L}_{i}(\mathcal{D}^{i}_{\textrm{TTT}};{\bm{\theta}}_{\textrm{i}})=\sum_{d\in\mathcal{D}^{i}_{\textrm{TTT}}}\left(\sum_{n=2}^{K}\mathcal{L}_{\textrm{LM}}(y_{n}|x_{1},y_{1},...,x_{n};{\bm{\theta}}_{\textrm{i}})+\mathcal{L}_{\textrm{LM}}(y_{\textrm{test}}|x_{1},y_{1},...,x_{K},y_{K},x_{\textrm{test}};{\bm{\theta}}_{\textrm{i}})\right)$ (6)
where $\mathcal{L}_{\textrm{LM}}$ is the standard cross-entropy loss for language modeling. Note that we include loss terms for demonstrations starting from the second example ($n=2$). By doing so, we encourage the model to start reasoning about the transformation pattern from the second demonstration pair itself.
Task-Specific Parameters:
Instead of learning a single LoRA adapter for all tasks in the test set, we learn an individual task-specific LoRA adapter for each task. That is, we obtain $N$ different LoRA adapters, where $N$ is the number of test tasks.
3.3 Results
Figure 3: Accuracy of different data and optimization ablations in TTT: Our data ablation studies reveal that the ICL data format is crucial for effective TTT, and that applying transformations to augment the TTT dataset notably enhances performance. In optimization ablations, learning task-specific adapters significantly outperforms using a single adapter. Additionally, taking a loss on the in-context demonstrations provides a minor performance boost, while using quantized LoRA results in only a slight performance decrease drop. Full discussion in Section 3.
We compare the main implementation of our method to the following ablations:
  1. 1.
    FT (No TTT): The vanilla baseline where TTT is ablated and the fine-tuned model is used instead.
  2. 2.
    No Transformations: No transformation-based data augmentation. That is, data from step 2 of the data generation pipeline described in Section 3.1 is not included in the test-time training dataset.
  3. 3.
    End-to-End (E2E) Data: Instead of the standard in-context task setup, we use the end-to-end task formulation, as described in Section 3.1.
  4. 4.
    Shared TTT: In contrast to learning a task-specific LoRA adapter, a single LoRA adapter is learned using an aggregated dataset of all tasks.
  5. 5.
    No Demonstration Loss: No loss is taken on the demonstrations in the training outputs of the data. That is, the TTT loss is simply:
    $\mathcal{L}_{i}(\mathcal{D}^{i}_{\textrm{TTT}};{\bm{\theta}}_{\textrm{i}})=\sum_{d\in\mathcal{D}^{i}_{\textrm{TTT}}}\left(\mathcal{L}_{\textrm{LM}}\left(y_{\textrm{test}}|x_{1},y_{1},...,x_{K},y_{K},x_{\textrm{test}};{\bm{\theta}}_{\textrm{i}}\right)\right)$ (7)
  6. 6.
    QLoRA: Rather than full-precision base model updates, quantized LoRA adapters (Dettmers et al., 2024) are learned for each task, which is the alternative for LoRA considered for memory efficiency.
[…]
Next, we ablate multiple components of TTT optimization to analyze their contribution to the performance. Learning a single LoRA adapter across all tasks reduces performance on 7 tasks ($\bf 24\%$ decrease). This is expected as learning a dedicated adapter allows more parameters to train per task. Second, the decision that we made by taking loss on the output demonstrations marginally improves the performance ($26\to 29$), as we believe that this forces the model to reason about the transformation while processing the demonstrations. Finally, we observe that using quantized LoRA (QLoRA) only leads to a marginal drop in performance ($29\to 26$) — in memory-bottlenecked scenarios using QLoRA may be viable.
4 What Inference Strategy After TTT?
Figure 4: Augmented inference and hierarchical voting (Section˜4): We use leave-one-out tasks and invertible geometric transformations to obtain multiple equivalent versions of the task for augmented inference. Predictions from these versions are aggregated with a hierarchical voting strategy: first, voting is performed within each transformation, and then the top candidates from each transformation undergo global voting to yield the top two predictions.
4.1 Augmented Inference
Recent work has shown that scaling test-time compute can significantly improve the performance of LMs. One of the most common techniques to do this is by sampling multiple responses, and then selecting the best response using a ranker. However, while sampling is very effective in domains with multiple possible solutions (programs in code) or multiple possible paths to the final answer (math), it can be detrimental when generating answers directly, as there is no way to directly enforce diversity across samples while ensuring coherence within samples. As an alternative inference-time scaling, we use an augmented inference strategy that generates multiple prediction candidates by using geometric transformations, combined with a greedy decoding scheme.
For a given task with training examples ${(x_{k},y_{k})}_{k=1}^{K}$ and test input $x_{\textrm{test}}$, we use invertible geometric transformations to produce equivalent transformed versions of the task, as shown in Fig.˜3. Let $\mathcal{T}$ be some set set of invertible geometric transformations (e.g., rotations and reflections). For each transformation $t\in\mathcal{T}$, we apply $t$ to all training demonstrations and the test input and run our model with these transformed inputs. We then apply the inverse transformation to obtain the final prediction for that transformation.
$\tilde{y}$ $\sim\textrm{LM}(t(\mathbf{d_{\textrm{input}}})):=\left[t(x_{1}),t(y_{1}),\dots,t(x_{\textrm{test}})\right]$ (8)
[…]
We further augment our predictions by permuting the order of training examples. For each transformation $g$, we sample $n=2$ different permutations of the demonstration sequence, resulting in $n\cdot|\mathcal{T}|$ total predictions per task. This is to mitigate any bias in the model’s processing of the demonstration sequence. Bober-Irizar and Banerjee (2024) also find transpose and rotation is helpful to produce extra prediction candidates.
4.2 Ensembling Predictions (Voting Strategy)
We employ a hierarchical voting strategy to determine the final prediction from the set of candidates $\{y\}_{i=1}^{n\cdot|\mathcal{T}|}$. This approach involves two stages of voting to progressively narrow down the best candidates: first, by selecting the most frequent predictions within each transformation, and then by conducting an overall vote across transformation-specific candidates to identify the top-2 most frequent predictions. The details of each stage are as follows:
  1. 1.
    Intra Transformation Voting: We group predictions by their corresponding transformation $t$ and select the top-3 most frequent predictions within each group. If fewer than 3 unique predictions exist within a group, we supplement the candidates by computing additional predictions through:
    • Row-based majority: For each row in the predicted output grid, we take the most frequent row values across all predictions in the transformation group.
    • Column-based majority: Similarly, for each column in the predicted output grid, we take the most frequent column values across all predictions in the transformation group.
  2. 2.
    Global Voting: Using the selected transformation-specific candidates obtained from (1), we conduct an overall vote to select the top-2 most frequent predictions for submission. In case of a tie, predictions with the identity transformation are given priority.
4.3 Results
Figure 5: Accuracy of different invertible transformations and voting schema: Our analysis shows that while individual transformations generally perform at a modest level and are comparable to one another, aggregating across them through voting yields substantial improvements. Notably, a hierarchical voting strategy with two voting stages surpasses a flat voting approach. Our hierarchical method approaches oracle-level performance, demonstrating its effectiveness in accurately selecting the correct answer when present. Full discussion in Section 4.3.
To analyze the impact of augmented inference and voting, we run the following ablations:
  1. 1.
    Vanilla: This baseline follows a standard inference approach without any augmented inference or voting. It generates 2 predictions from the model for $2$ permutations of the task. This setup serves as a reference point to assess the benefits of our augmented inference and voting strategy.
  2. 2.
    Transformed Inference (Rotate/Transpose/Flip): Measures performance when predictions are generated solely from a specific transformed version of the task shown in Fig.˜5. This assesses the individual effectiveness of each transformation applied in isolation. Note that Vanilla can also be considered a part of this category, with the transformation being the identity function .
  3. 3.
    Hierarchical Voting: Our full pipeline, which includes both augmented inference and voting.
  4. 4.
    Flattened Voting: Instead of using a hierarchical voting strategy, we perform a single voting round on the full set of $n\cdot|\mathcal{T}|$ predictions to identify the $2$ most frequent predictions.
  5. 5.
    Oracle: The oracle selects the correct answer if it exists in the set of $n\cdot|\mathcal{T}|$ predictions. The oracle provides an upper-bound on the best performance possible if the voting procedure was perfect.
The results are summarized in Figure 5. As shown in the figure, the individual performance of specific transformed versions is generally poor, with the transpose transformation yielding the worst accuracy. However, aggregating across these transformations through voting procedures leads to significant improvements. This suggests that some tasks may be easier to solve in their transformed versions, and that using self-consistency (voting) for aggregation is generally beneficial, a finding also observed in prior work. Additionally, while the flattened voting procedure improves accuracy, our hierarchical voting procedure outperforms it. In fact, our hierarchical procedure is comparable to the oracle, indicating that hierarchical aggregation effectively selects the correct answer (when it exists) with high accuracy.
5 What Fine-Tuning Before TTT?
Figure 6: LLM based synthetic tasks generation: Given some seed task descriptions and task generator functions in Python, we generate more generator functions to produce novel tasks. We use three different approaches: (1) few-shot prompting with only generators, (2) few-shot prompting with generators and task descriptions, (3) two-stage approach: first generate free form descriptions, then condition on them to generate more generators (shown in Fig.˜9).
While test-time training facilitates task-specific adaptation, the base model’s capabilities impacts the final performance. We developed several approaches for generating synthetic training data to enhance the base model’s abstract reasoning capabilities through fine-tuning, exploring both automated and semi-automated methods for task generation. In this section, we detail our fine-tuning data generation strategies and analyze the impact of different data sources and model sizes on final performance.
5.1 Preparing Fine-tuning Data
Hodel (2024) provides domain-specific language (DSL), ReARC, as well as the transformation $f_{i}$ that solves the task-$i$, and the data generation function $g_{i}$ that are implemented in this DSL for each training task in the $\mathcal{D_{\textrm{ARC}}^{\textrm{train}}}$ dataset. These functions enable sampling of new input-output pairs that maintains the same underlying transformation principle:
$d=(x,y)\sim\textrm{eval}(g_{i})$ (10)
where $d$ represents a newly generated input-output pair that can be solved using the same transformation function $f_{i}$ as the original task-ii [^2] We can verify the generated examples by asserting $f_{i}(x)=y$..
(a) Using Existing Generators
The generator functions $g$s in ReARC already provide an effective data augmentation tool by producing different instantiations of same tasks. We generate extra samples from these training tasks by running these codes many times and randomly splitting these new examples ($d\sim\textrm{eval}(g_{i})$) to set of train and test examples. These augmented examples are already provided with their DSL release.
(b) Few-shot Prompting an LLM
Additionally, we used several approaches to generate novel tasks using an LM (in our case an ensemble of GPT4 and GPT4-o).
The simplest approach generates new task generators using few-shot examples:
$g^{\prime}\sim\textrm{LM}(g_{1},g_{2},\dots,g_{m})$ (11)
where $g^{\prime}$ is a new generator function and $g_{1},\dots,g_{m}$} are existing generator functions (shown in Fig.˜6)s. We sample different $m$ examples by uniformly from existing training set. We repeat this process multiple times to get a good amount of tasks.
We augment the generator functions with task descriptions and jointly generate both descriptions and generators:
$(s^{\prime},g^{\prime})\sim\textrm{LM}({s_{1},g_{1},s_{2},g_{2},\dots s_{m},g_{m}})$ (12)
where $s_{i}$ represents the description of task $i$.
To get the task descriptions, we manually created seed descriptions for 10 training tasks. These seed descriptions were then used to generate descriptions for the training and validation tasks through few-shot prompting. To increase diversity of tasks we use task descriptions with hierarchical fields (category, summary, and description). The process of getting these descriptions provided in the Section˜D.1.
Instead of jointly generating task descriptions and function generations, we additionally deployed a two-stage approach described as following:
$s^{\prime}$ $\sim\textrm{LM}(s_{1},s_{2},\dots s_{m})$ (13)
$g^{\prime}$ $\sim\textrm{LM}(s_{1},g_{1},s_{2},g_{2},\dots,s_{m},g_{m},s^{\prime})$ (14)
This approach first generates a task description $s^{\prime}$ and then conditions the generator creation on both existing task pairs and the new description. In total we collected 6426 generators with these LLM based approaches. We provide qualitative samples from these LM generated tasks in Fig.˜11
(c) Geometric Transformations
Finally, our synthetic tasks are enhanced through various geometric transformations, such as basic transformations (rotations, reflections, random shift and size scaling), pattern operations (random patching, tiling, and repetition), color permutations, and composite transformations involving sequential application of multiple basic transformations. These transformations are applied in three ways:
  • Input grids only: $(x,y)\rightarrow(t(x),y)$
  • Output grids only: $(x,y)\rightarrow(x,t(y))$
  • Both input and output: $(x,y)\rightarrow(t(x),t(y))$
The complete specification of transformations and their application details are provided in Section˜B.1. These transformations are applied randomly to variants of tasks with 30% of the time.
5.2 Results
We perform full fine-tuning 1B, 3B Llama 3.2 instruction-tuned, and 8B Llama 3 instruction-tuned using augmented data. The format and training objective is same as the ones described for TTT in Section˜2.4. Hyper-parameter details are given in Section˜B.2. We do the following ablations for augmented data:
  1. 1.
    No FT: The original Llama 3 instruction-tuned model without any fine-tuning.
  2. 2.
    All: We use all methods described in Section 5.1, including ReARC, rule-based augmentation, and LM generation.
  3. 3.
    No-Geom: We remove geometric transformations from all tasks.
  4. 4.
    No-LM: We only use ReARC and rule-based augmentation, excluding tasks generated by the LM.
Figure 7: Left: Accuracy when fine-tuning with different data sources. While all fine-tuned models perform similarly, their performance after TTT shows considerable variance. As expected, removing geometric transformations from the fine-tuning data reduces performance compared to the model trained on the full dataset. Surprisingly, excluding LM-generated data from fine-tuning actually outperforms the model trained on all data. Right: Performance results across different model sizes. As expected, performance of the base fine-tuned model improves with increasing model size, aligning with current scaling law trends. However, the scaling behavior after TTT is less clear. For instance, the final performance of the 1B and 3B models is identical after TTT. Full discussion in Section 5.2.
How does FT data affect TTT?
We compare models using different fine-tuning data in Fig.˜7. We find that the model trained on ReARC with rule-based augmentation achieves the strongest performance. Surprisingly, including LM-generated tasks hurts performance by 5%, indicating that current LM-based task generation methods may need more sophisticated filtering mechanisms as used in Li et al. (2024) (see their results in Table˜1). Finally, we find that FT performance shows little correlation with TTT performance.
Model Size and Scaling in TTT
We show results using different model sizes in Fig.˜7. Increasing the model size consistently improves FT performance, with the 8B model achieving the highest accuracy of 36%. We also observe that TTT effectively closes the performance gap for smaller models, with the 1B and 3B models achieving similar accuracy after TTT.
6 ARC Benchmark and Comparison to Other Systems
Table 1: Scores of different systems on the ARC validation set: Our TTT pipeline improves base models consistently. We achieve 47.1% accuracy when applied to our fine-tuned model, 53% when applied to BARC model from Li et al. (2024), achieving state-of-the-art on pure LM based approaches. We ensemble our method with program synthesis based models, where we achieve (61.87561.875%) state-of-the-art performance comparable to average human performance (60.2%).
Program Synthesizer Fine-tuned LM TTT Method Score (pass@2)
X Ours X 18.2518.25%
X Ours Ours 47.12547.125%
X BARC Ours 5353%
BARC Ours Ours 58.558.5%
BARC BARC Ours 61.87561.875%
Avg. Human 60.260.2%
Best Human 97.897.8%
BARC (ensemble) 54.37554.375%
BARC (no synthesizer) 39.2539.25%
Claude - Few-shot prompting 2121%
GPT-4.0 - Few-shot prompting 99%
Following our development experiments on 80 tasks, we present comprehensive results on the full ARC public evaluation set, comparing our system against existing approaches. Our analysis focuses on three key aspects: the impact of our TTT methodology, the benefits of combining our approach with existing methods and the differences between fully neural and program synthesis methods.
Impact of Test Time Training
We applied to our TTT and inference procedure (explained in Section˜3 and Section˜4) to our base fine-tuned models (fine-tuned 8B model without any LM data in Section˜5). TTT improves accuracy from 39.2539.25% to 47.12547.125%, surpassing existing end-to-end neural model results.
Integration with Existing Methods
A concurrent work by Li et al. (2024) introduced BARC, achieving 54.37554.375%accuracy by combining neural and program synthesis approaches—previously the highest publicly available result. While their fully neural approach shares similarities with our system, our TTT and inference pipeline has several additional components that boost performance. In particular, our test-time-training includes per-task LoRA and a larger set of augmentations, while our prediction pipeline includes an augmented inference under invertible transformations and a hierarchical self-consistency voting scheme. To validate our improvements, we applied our TTT pipeline to BARC’s fully neural model, achieving 53% accuracy—a 35% improvement over their original TTT method.
Building on these results, we explored various combinations of our approach with BARC’s components:
  • Combining our TTT pipeline and neural model with BARC’s synthesizer raised accuracy to 58.558.5%.
  • Combining our TTT pipeline with BARC’s neural model and synthesizer raised accuracy to 61.87561.875%.
This final configuration establishes a new state-of-the-art on the ARC public evaluation set, matching the average human performance (LeGris et al., 2024b). While this represents significant progress, there remains a substantial gap to the best human performance of 97.8%, indicating room for further improvements.
Comparing Program Generation and End-to-End Modeling
Li et al. (2024) found that program synthesis and fully neural predictors for ARC are highly complementary, even when trained on the same tasks. Their end-to-end neural model can only solve 42.2% of the tasks solved by the program synthesis model. However, we find that when equipped with our TTT pipeline, BARC’s fine-tuned fully neural model solves 73.5% of the tasks that are solved by the program synthesis model. This suggests that our TTT pipeline significantly improves the neural model’s ability to learn systematic reasoning patterns similar to those captured by program synthesis models.
7 Conclusion
In this work, we conduct an investigation of test-time training and demonstrate that it can significantly improve LM performance on the popular ARC dataset. We find that learning task-specific LoRA adapters and generating augmented test-time datasets using geometric transformations are crucial. We also develop an augmented inference pipeline that uses invertible transformations to generate multiple predictions and then uses self-consistency to select the best candidates. Our overall pipeline applies multiple test-time computation methods, with each component contributing positively. This suggests that not only can test-time compute improve LM performance, but different test-time methods can also complement one another. Our TTT pipeline, combined with an existing method (BARC), achieves state-of-the-art results on the ARC public set and performs comparably to an average human. Our findings suggest that test-time methods could play a pivotal role in advancing the next generation of LMs.
Limitations
Evaluation Framework
[…]
Experimental Reproducibility
Given the computational requirements of our experiments, this preprint reports results without comprehensive standard error analysis. Our preliminary observations indicate minimal variance across runs, and we plan to include detailed statistical analysis in the final version.
Data Leakage
Even though the base Llama-3 perform extremely poorly on the public validation set, the public availability of the dataset on various platforms (GitHub, Kaggle) introduces the possibility that these models may have encountered these examples during pre-training.
Acknowledgments
We thank Aniruddha Nrusimha for helpful discussions on parameter efficient training, and Jyo Pari for feedback on early drafts of this paper. This work was supported by the National Science Foundation under grants IIS-2212310, IIS-2238240, and CCF-2217064.
Appendix A ARC Dataset
We present the tasks in development set, format and evaliation for the ARC dataset (available in this https link.).
A.1 Data Format
We use numpy’s array printing format for all the experiments as shown in Fig.˜8.
A.2 List of 80 Tasks Used For Development
We use following (Table˜2) tasks validation tasks for our development.
Table 2: Selected development tasks and their hardness level based on LeGris et al. [2024b].
ID Level ID Level ID Level ID Level
0a1d4ef5 easy 762cd429 medium e5c44e8f hard e99362f0 expert
692cd3b6 easy e7639916 medium 604001fa hard 1acc24af expert
1da012fc easy e1d2900e medium 4364c1c4 hard f9a67cb5 expert
66e6c45b easy aee291af medium 506d28a5 hard ad7e01d0 expert
3194b014 easy e95e3d8e medium 2037f2c7 hard ea9794b1 expert
963f59bc easy e0fb7511 medium d5c634a2 hard 58e15b12 expert
d37a1ef5 easy ae58858e medium ac605cbb hard 891232d6 expert
358ba94e easy 93c31fbe medium 27f8ce4f hard 5833af48 expert
f3cdc58f easy 27a77e38 medium 66f2d22f hard 4ff4c9da expert
55059096 easy 9bebae7a medium 3ed85e70 hard 5b692c0f expert
c7d4e6ad easy 9ddd00f0 medium 8b28cd80 hard e2092e0c expert
4b6b68e5 easy fe9372f3 medium d19f7514 hard 47996f11 expert
00576224 easy 69889d6e medium dc2aa30b hard 34b99a2b expert
a04b2602 easy 15663ba9 medium f5c89df1 hard 1c56ad9f expert
e9c9d9a1 easy 17b80ad2 medium 50f325b5 hard e6de6e8f expert
ef26cbf6 easy 16b78196 medium 08573cc6 hard fea12743 expert
7ee1c6ea easy 5b6cbef5 medium 3d31c5b3 hard 31d5ba1a expert
e9ac8c9e easy 40f6cd08 medium 94133066 hard 79fb03f4 expert
1a2e2828 easy 505fff84 medium 136b0064 hard 8719f442 expert
770cc55f easy d017b73f medium 90347967 hard a8610ef7 expert
A.3 Evaluation
We follow the competition rules and in any of the two pass@2 predictions of the system is correct, we consider that test as correct. In the reported task level accuracies, we did not give partial points if all tests are not solved, except the final table Table˜1.
Figure 8: Data Format: We convert grids to strings by representing them as numpy arrays of digits from 0 to 10 where each digit corresponds to a different color.
Appendix B TTT
We present transformation used in TTT and the training details.
B.1 Transformations
We provide the augmentations used in TTT in the Section˜3, please refer to our code base for their implementations. After application of these augmentation, we additionaly shuffle colors and shuffle training examples. Note that these transformations are applied to all input and output grids.
Table 3: We provide the augmentations use in our TTT procedure with their function signature and description.
Augmentation Name Description
Rotate(90) Rotates a grid 90 degrees.
Rotate(270) Rotates a grid -90 degrees.
Rotate(180) Rotates a grid 180 degrees.
Flip(0) Flips a grid horizontally
Flip(1) Flips a grid vertically
Reflect(0, reverse=True) Flips a grid horizontally and prepend to the left of the original grid
Reflect(1, reverse=True) Flips a grid vertically and prepend to the above of the original grid
Reflect(0, reverse=False) Flips a grid horizontally and append to the right of the original grid
Reflect(1, reverse=False) Flips a grid vertical and append to the left of the original grid
RandomTranslateXY() Shifts a grid randomly both in horizontal and vertical directions. The maximum shift size is 4
Transpose() Reflect a grid on diagonal
IncreaseResolution(2) Upscale the grid by interleaving elements in both horizontal and vertical directions
IncreaseHeight(2) Upscale the grid by interleaving elements in vertical direction
IncreaseWidth(2) Upscale the grid by interleaving elements in horizontal direction
Chain([Rotate(90),IncreaseResolution(2)]) Sequential application of Rotate(90) and IncreaseResolution(2)
Chain([Rotate(270),IncreaseResolution(2)]) Sequential application of Rotate(270) and IncreaseResolution(2)
Chain([Rotate(180),IncreaseResolution(2)]) Sequential application of Rotate(180) and IncreaseResolution(2)
Chain([Flip(0),IncreaseResolution(2)]) Sequential application of Rotate(180) and IncreaseResolution(2)
Chain([Flip(1),IncreaseResolution(2)]) Sequential application of Rotate(180) and IncreaseResolution(2)
Chain([Transpose(),IncreaseResolution(2)]) Sequential application of Rotate(180) and IncreaseResolution(2)
B.2 Training Setup & Hyperparameters
Table 4: TTT hyper-parameters
Hyper parameter Search Space
$r$ lora rank [128]
$\alpha$ lora alpha 16
learning rate [5e-5, 1e-4]
epochs 2
batch size [1, 2]
optimizer AdamW [Loshchilov and Hutter, 2018]
We use torchtune[torchtune maintainers and contributors, 2024] library to train LoRA adapters on Llama-3 family of models. We apply LoRA training to query and value projection weights of the self-attention layer, to the MLP weights and to the output projection layer (was only available for Llama-3 8B in torchtune). We present hyper-parameters of this training in Table˜4.
Appendix C Inference
We resort to vLLM [Kwon et al., 2023] library for prediction as it provides fast kernels and batched inference for our models and LoRA inference. We just use greed decoding as we did not see improvements with temperature sampling in our early experiments. We use 90, 180 degree rotations, horizontal, vertical, and diagonal (transpose) flips as our invertible transformations.
Appendix D LM Data Generation
We described three approaches in Section˜5 to use LM, we generated 6426 task generators by few-shot prompting GPT-4 and GPT-4o models [OpenAI, 2023, Hurst et al., 2024].
D.1 Getting Descriptions for Tasks
This procedure is shown in Fig.˜10. We initially described 10 training tasks with the hierarchical-style shown in Fig.˜6. Then, for other training tasks tasks, we obtained less quality crowd-worker annotations from LARC [Acquaviva et al., 2022] project. By using our high-quality seed annotations and their LARC version, we 10-shot prompt and LM to produce high quality annotations for the other training tasks.
You are an intelligent agent that can induce task descriptions from examples. For Category, please *do not* use generic terms like Transformation, Pattern Recognition. —————- Task: {stringified task inputs and outputs} LARC Description: {description of the task-1 from LARC dataset} Good Description: {hierarchical description} —————- $[truncated]$ —————- Task: {stringified task inputs and outputs for task-K} LARC Description: {description of the task-K from LARC dataset} Good Description: {hierarchical description} —————- Task: {stringified task inputs and outputs for query task} LARC Description: {description of the query task from LARC dataset}
D.2 Few-shot Prompting Details
We use the following simple prompting template with k-shot prompting for all data generation procedures, where numbers filled with examples sampled from seed set. In simple few-shot generation, we exclude examples. We use GPT-4 and GPT-4o to generate the new scripts.
You are a problem generator on 2D grids of colors. Here are some examples of such transformations, please follow the format: —————- Example: {description of the generator function-1} Script: {generator function-1} —————- $[truncated]$ —————- Example: {description of the generator function-K} Script: {generator function-K} Please generate more and make sure they are different:
Figure 9: Two-stage generation using an LLM: First, we prompt the LLM to generate a task description using few-shot prompting. Then, we generate the new generator based on existing task pairs and the newly created description.
Figure 10: Generating quality seed descriptions: We use few-shot prompting to generate descriptions for a given task, using 10 manually created seed descriptions along with crowd-worker annotations from Acquaviva et al. [2022] as few-shot examples. For a given new task, we similarly provide the LM with examples and crowd-worker annotations (available only for training tasks).
Appendix E Fine-tuning
In each dataset described in Section˜5, we generated approximately $600000$ ARC tasks from the available task generator functions (e.g. training tasks, ReARC tasks, and LM generated tasks) by repeatedly picking 2-7 examples from the generated input outputs, and also applying geometric transformations if used.
E.1 Fine-tuning Transformations
We use all the transformations given in Table˜3, and some additional transformations given in Table˜5. In the fine-tuning case, different from TTT, we apply augmentations to only inputs, only outputs or both.
Table 5: We provide the additional augmentations use in our data generation for fine-tuning with their function signature and description.
Augmentation Name Description
Repeat(direction, n) Rotates a grid in horizontal or vertical direction by $n$ times.
DropoutOutput Randomly deletes some patches of the output grids.
DropoutInput Randomly deletes some patches of the input grids
E.2 Fine-tuning Hyper-parameters
We perform full fine-tuning on LLama-3 family models by using torchtune library. We train each model up to 16000 steps. We use 2xNVIDIA A100 GPU for 1B models, 4xNVIDIA A100 GPU for 3B and 8B models. We present hyper-parameters in Table˜6.
Table 6: Fine-tuning hyper-parameters
Hyper parameter Search Space
learning rate 2.5e-5
epochs 2
batch size 32
optimizer AdamW [Loshchilov and Hutter, 2018]
scheduler Cosine LR Schedule with 2k warmup
E.3 Qualitative Examples
Figure 11: Example tasks generated by LM data augmentation procedure: We display three reasonable tasks that we can infer a simple transformation (valid), and three tasks that we could not infer a simple transformation (invalid).
We present some qualitative examples from our LLM data generation procedure in Fig.˜11.